OpenStudy (anonymous):

Write an equation for the line perpendicular to y = 5x – 12 that contains (–3, –1)

OpenStudy (anonymous):

use y=mx+b :)

OpenStudy (anonymous):

so lets take a look at this... what do you know about the slope of perpendicular lines?

OpenStudy (johnweldon1993):

Perpendicular...means it has the negative reciprocal slope of the other line's slope...what is the negative reciprocal of 5...?

OpenStudy (johnweldon1993):

I'll let @iheartfood continue :)

OpenStudy (anonymous):

hahah kk @johnweldon1993 feel free to jump in tho if u want! :) but that's right! perpendicular means that it will have a slope that is opposite reciprocal :)

OpenStudy (anonymous):

noting really i just got back in to geometry i didnt really learn a whole lot

OpenStudy (anonymous):

okay... so basically that's all you need though! :) correct me if i'm wrong @johnweldon1993... i might be forgetting something here lol... but i do believe that you are solving for the y-int by plugging in. so -1=m(-3) + b.. solving for b but @johnweldon1993 i'm forgetting, are we solving with 5 as the m value? or -5?

OpenStudy (anonymous):

so it would be y+1=-1/5(x+3)

OpenStudy (anonymous):

no, not quite.. @johnweldon1993 can you please take a look when you get the chance? @redcarter we are solving for "b" ur y-int using either -1=(5)(-3) + b or -1=(-5)(-3) +b but i can't remember which.. so lets wait so he/she can take a look :)

OpenStudy (anonymous):

ok

OpenStudy (primeralph):